Kwame’s age is six(6) times that of his brother. In Ten(10) years time, kwame will be four times as old his brother. What are their present ages.?

Answer by ikleyn(52790)

.

Let x be the brother's age now.


Then the Kwame's age is 6x years.


In 10 years


    4*(x+10) = 6x + 10.


Simplify and solve for x


    4x + 40 = 6x + 10

    40 - 10 = 6x - 4x

    30      = 2x

     x      = 30/2 = 15.


ANSWER.  Their present ages are 15 and 90 years.


         If you tell me that it is impossible, I will answer you that it may happen in parallel Universe.